OpenStudy (anonymous):

how many gallons of elastomeric paint are required to paint the exterior surface area of an open cylindrical concrete tank of height 3.6m and base outside diameter of 2.0m if one gallon of paint covers 10m^2

OpenStudy (anonymous):

First let us find out the area of the cylindrical tank where paint has to be applied: so, Area of the tank = Curved surface area + one base are \[ \huge = 2\pi r h+ \pi r^2 = \pi r (2h+r)\] We are give that: h = 3.6 m and r = d/2= 2/2 = 1 m Substituting these values in the above eq we have. \[ \huge = 3.14 \times 1 \times (2\times 3.6+1)\] \[ \huge = 3.14 \times (7.2+1) = 3.14 \times 8.2 =25.748 m^2\] Amount of elastomeric paint required = (area of the cylindrical tank) /10 m^2 \[\huge = \frac{25.748}{10} = 2.5748 \approx 2.6 \] Hence approximately 2.6 gallons of elastomeric paint would be required to paint the cylindrical tank. @shainarobles

OpenStudy (anonymous):

@shainarobles Are you there?

OpenStudy (anonymous):

yes, thankyou.